OpenStudy (anonymous):

State whether every point on the graph is a solution of the equation. Explain why or why not. x – 4y = –4 A. Every point shown on the graph is not a solution of the equation because when the coordinates of the points are substituted into the equation, the equation is not true for all the points. B. Every point shown on the graph is a solution of the equation because when the coordinates of the points are substituted into the equation, the equation is true for all the points.

OpenStudy (anonymous):

i give medals

OpenStudy (igreen):

Plug the three points into the equation and see if they are true. (-4, 0), (0, 1), (3, 2) x - 4y = -4 (-4, 0) -4 - 4(0) = -4 x - 4y = -4 (0, 1) 0 - 4(1) = -4 x - 4y = -4 (3, 2) 3 - 4(2) = -4

OpenStudy (igreen):

Simplify those 3 equations and see if they are true.